# Dual MPC for Active Learning of Nonparametric Uncertainties

### v1

@@ IV-B Safety Guarantees
where $\lambda\in[0,1]$ is a parameter that weights the contingency horizon (7b) and the performance horizon (7c). This MPC uses a secondary robust input sequence $\bar{U}_{k}$, based on the RMPC from (3), to ensure the safety of the GP-MPC. By learning the dynamics $g$, the performance horizon can select a ‘better’ control input $\hat{u}^{*}_{0\mid k}$, while the contingency horizon restricts the input $\hat{u}^{*}_{0\mid k}=\bar{u}^{*}_{0\mid k}$ to the robust admissible input set. Consequently, the contingency MPC can reduce the conservatism of the RMPC while ensuring safety [geurts_baltussen_2025].
@@
Let $\mathbb{D}_{i\mid k}:=\{d(\mathbf{z}_{i\mid k},\mathcal{I}_{k})\mid\mathbf{z}_{i\mid k}\in\mathbb{X}\times\mathbb{U}\}$ denote the set of all possible GP mean predictions at prediction step $i$ given the data $\mathcal{I}_{k}$ and $\hat{\mathbb{X}}_{i\mid k}:=\{\hat{x}_{i\mid k}\in\mathbb{X}\ominus 2\sigma_{x,i\mid k}\}$ the tightened state constraints of the performance horizon. The learned residual dynamics satisfy $\mathbb{D}_{i\mid k}\in\mathbb{W}$, and it holds that $\hat{\mathbb{X}}_{i\mid k}\supseteq\bar{\mathbb{X}}_{i}$, for $i=0,1,\dots,N$ and for all $k\in\mathbb{N}$.
@@ V-A Active Learning Framework
In this paper, we extend our work in [baltussen2025] to leverage the GP-MPC for the active learning of nonparametric uncertainties. We extend the active learning framework from [soloperto_augmenting_2020] as it provides an intuitive framework to augment existing MPC formulations with active learning. By combining this with the multi-horizon contingency MPC, we can leverage the dual effect of the performance horizon to actively learn the model while the contingency horizon ensures safety. This framework uses a learning cost function $H$ while the primary cost function $J$ appears in additional constraints.
@@
Here, the primary cost function $J$ of the performance horizon is composed of a baseline cost $J_{k}^{B}$ and a deterioration term $\Delta_{k}$, which quantifies the cost incurred for learning. The deterioration used for exploration $(\Delta_{k})$ is bounded over time by (9c) and at each time step by (9d). A storage variable $Y_{k}:=Y_{k-1}+\bar{\beta}\,\max(J^{+}_{k},0)+\bar{\gamma}-\Delta_{k}$ accumulates cost deterioration, preventing excessive performance loss over time. The term $J^{+}_{k}:=J(x_{k-1},\tilde{x}_{0\mid k-1},\tilde{U}_{k-1})-J^{B}_{k}$, where $\tilde{x}_{0\mid k-1},\tilde{U}_{k-1}$ denote the previous solution of (9), allows exploration when it can potentially improve performance [soloperto_augmenting_2020]. The learning behavior is tuned via the user-defined parameters $\bar{\beta}$, $\bar{\gamma}$, $\beta^{\text{max}}$, and $\gamma^{\text{max}}$, see [soloperto_augmenting_2020] for details. We define the baseline cost $J_{k}^{B}$ as the optimal cost $J(x_{k},\hat{x}^{*}_{0\mid k},\hat{U}^{*}_{0\mid k})$ from the performance horizon of (7), which requires solving both (7) and (9) at each time step. Alternatively, an upper bound on $J^{B}_{k}$ can be computed in different ways (see [soloperto_augmenting_2020]).
@@
$H(x_{k},\hat{x}_{0\mid k},\hat{U}_{k})=-\sum_{i=0}^{N}\Sigma^{d}(\mathbf{z}_{i\mid k},\mathcal{I}_{k}).$ (10)
@@ V-B Safety Guarantees
Similar to [soloperto_augmenting_2020], the robust recursive feasiblity of the active learning dual MPC is ensured by relying on a robust MPC formulation. In contrast to [soloperto_augmenting_2020], we only require the constraints to be robustly enforced for the contingency horizon [geurts_baltussen_2025] which can reduce conservatism. The robust recursive feasibility of the contingency MPC (7) is then ensured by using a robust MPC in the contingency horizon [geurts_baltussen_2025].
@@ VI Numerical Example
We perform a numerical study on a nonlinear mechanical system to compare the discussed MPC algorithms and demonstrate the proposed active learning dual MPC. Specifically, we consider a mass–spring–damper system, shown in Fig. 1, with an unknown nonparametric spring force $g_{c}(x,u)$, that is described by the continuous-time state-space equation:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\underbrace{\begin{bmatrix}x_{2}\\ -\frac{1}{m}\left(c\,x_{2}+k\,x_{1}+u\right)\end{bmatrix}}_{=\,f_{c}(x,u)}+\underbrace{\begin{bmatrix}0\\ -\frac{1}{m}k(e^{-x_{1}}-1)x_{1}\end{bmatrix}}_{=\,Bg_{c}(x,u)},$
where $x_{1}$ [m] is the position, $x_{2}$ [m/s] is the velocity, $m=1$ [kg] is the mass, $c=1$ [Ns/m] is the damping coefficient and $k=0.33$ [N/m] is the spring constant. We consider $v_{k}=0$. The system is subject to state constraints $\mathbb{X}=[-0.1,1.1]\times[-5,5]$, and input constraints $\mathbb{U}=[-5,5]$. The goal is to control the system to a desired position $x_{\text{ref}}^{\top}=\begin{bmatrix}1&0\end{bmatrix}$ and $x_{\text{ref}}^{\top}=\begin{bmatrix}1.095&0\end{bmatrix}$ after $t=5$ s while actively learning the unknown nonlinear spring force using the proposed dual MPC. The continuous-time system is linearized at the origin and discretized using the forward Euler method with a sampling time of $0.1$ seconds, yielding the nominal discrete-time model $f$. The input is applied with a zero-order hold and is kept constant over the sampling period. The primary cost function is a standard tracking cost
@@
VI-A MPC Implementations
We consider the primary cost function (11) with weighting matrices $Q=\text{diag}(20,1)$ and $R=1$ with $P$ from the LQR corresponding to $Q$ and $R$, and $\lambda=10^{-3}$. We consider a prediction horizon of $N=20$. For the GP regression, we use $x_{1}$ as the regression feature with a squared exponential kernel $k$ with a signal variance $\sigma_{f}^{2}=0.33$, noise variance $\sigma_{w}^{2}=0$ and length scale $l=0.3$. We use $M=4$ inducing points. The unmodeled spring force only affects the velocity dynamics, hence, $B=[0,1]^{\top}$. We consider $\kappa\equiv 0$ and $w_{k}\in\mathbb{W}$ for all $k\in\mathbb{N}$ with $\mathbb{W}=\{0\}\times[-0.33,0.33]$. The parameters of the active learning framework are manually tuned and set to $\bar{\beta}=1$, $\bar{\gamma}=0$, $\beta^{\text{max}}=0$, and $\gamma^{\text{max}}=\infty$, so that the degree of learning decays over time as the tracking cost decreases. The MPC parameters are constant for all MPC formulations.
The RMPC (3) robustly accounts for all possible realizations of the unmodeled dynamics and enforces robustly tightened constraints with an RCI terminal set, computed using the LQR, to ensure recursive feasibility. The GP-MPCs leverage the online collected measurements to identify and predict the unmodeled dynamics and use adaptive constraint tightening based on the predicted state covariance. Furthermore, the GP-MPCs uses the RMPC (3) in the contingency horizon. The passive GP-MPC [geurts_baltussen_2025] only leverages the cautious property of the dual effect. The proposed dual GP-MPC (9) uses the learning objective (10) to leverages both the cautious and explorative properties of the dual effect (Section III-A). Finally, we evaluate the active learning framework by [soloperto_augmenting_2020] that does not use a secondary horizon and relies on robustly tightened constraints from (3), and, therefore, lacks the cautious control property of the dual effect.
@@ VI-B Results
The results of the simulation study are shown in Fig. 2. The RMPC considers the worst-case disturbance caused by the uncertain spring stiffness. Initially, the RMPC overshoots and has a large steady-state error (2.74%). After $t=5$ s, the improved tracking is largely incidental. As the nominal model is unaware of the reduced stiffness, the controller tends to apply large control inputs, and the system happens to settle near the constraint, yielding a small steady-state error (0.48%) at $t=10$ s. The active learning MPC with a single horizon [soloperto_augmenting_2020], actively explores the state space and later exploits the learned model. Although the learned model is aware of the reduced stiffness, it is still constrained to the robust constraints that were designed for the worst-case disturbance. Consequently, the MPC cannot fully exploit the learned model, leading to conservative behavior and steady-state errors of 3.16% and 3.53%, at $t=5,10$ s, respectively.
The passive GP-MPC with the contingency horizon [geurts_baltussen_2025] is able to successfully control the system to the desired setpoint while learning and adapting to the nonparametric unmodeled dynamics. As the two horizons branch after the first input, the contingency horizon ensures that the RCI set is reachable at all times, while the performance horizon can exploit the learned model to reduce conservatism, and improve performance, significantly reducing the steady-state errors to 0.17% and 0.13%, at $t=5,10$ s, respectively. Lastly, the proposed active learning GP-MPC can safely excite the system to learn the unmodeled dynamics, and exploit the learned model to obtain similar performance (0.19% and 0.13%). The active learning GP-MPC (9) has an average CPU time of $0.2951$ seconds on a standard laptop PC using MATLAB, CasADi [andersson_casadi_2019] and IPOPT [wachter_implementation_2006], demonstrating the practical computational efficiency of the proposed approach.
@@ VII Conclusion
In this work, we proposed a dual MPC that leverages the dual effect of our Gaussian Process (GP) MPC to actively learn nonparametric uncertainties. We extended the active learning framework of [soloperto_augmenting_2020] by including the dual effect and adopted a multi-horizon contingency MPC formulation [geurts_baltussen_2025] to ensure robust constraint satisfaction. The resulting dual MPC features caution and exploration, the two properties of the dual effect, enabling safe control and active learning near constraints, and was demonstrated on a nonlinear system.
Although a simple problem was considered for the sake of analysis, the results show that both passive and active learning GP-MPCs with a contingency horizon can safely identify and control systems with unmodeled dynamics. Furthermore, the results indicate that the contingency horizon can reduce conservatism compared to single-horizon robust MPCs. The active learning GP-MPC is expected to outperform the passive variant in more complex systems that require exploration to effectively learn the unknown dynamics. Future work will focus on the performance analysis of both active and passive learning settings of the contingency MPC and explore the use of robust adaptive MPC to further reduce conservatism.

### v2

@@ IV-B Safety Guarantees
where $\lambda\in[0,1]$ is a parameter that weights the contingency horizon (7b) and the performance horizon (7c). This MPC uses a secondary robust input sequence $\bar{U}_{k}$, based on the RMPC from (3), to ensure the safety of the GP-MPC. By learning the dynamics $g$, the performance horizon can select a ‘better’ control input $\hat{u}^{*}_{0\mid k}$, while the contingency horizon restricts the input $\hat{u}^{*}_{0\mid k}=\bar{u}^{*}_{0\mid k}$ to the robustly admissible input set. Consequently, the contingency MPC can reduce the conservatism of the RMPC while ensuring safety [geurts_baltussen_2025].
@@
Let $\mathbb{D}_{i\mid k}:=\{d(\mathbf{z}_{i\mid k},\mathcal{I}_{k})\mid\mathbf{z}_{i\mid k}\in\mathbb{X}\times\mathbb{U}\}$ denote the set of all possible GP mean predictions at prediction step $i$ given the data $\mathcal{I}_{k}$ and $\hat{\mathbb{X}}_{i\mid k}:=\{\hat{x}_{i\mid k}\in\mathbb{X}\ominus 2\sigma_{x,i\mid k}\}$ the tightened state constraints of the performance horizon. The learned residual dynamics satisfy $\mathbb{D}_{i\mid k}\subseteq\mathbb{W}$, and it holds that $\hat{\mathbb{X}}_{i\mid k}\supseteq\bar{\mathbb{X}}_{i}$, for $i=0,1,\dots,N$ and for all $k\in\mathbb{N}$.
@@ V-A Active Learning Framework
We extend the passive GP-MPC (7) to leverage the posterior GP covariance for the active learning of nonparametric uncertainties. To this end, we extend the active learning framework from [soloperto_augmenting_2020] as it provides an intuitive framework to augment existing MPCs with active learning. By combining this with the multi-horizon contingency MPC, we can leverage the dual effect of the performance horizon to actively learn the model while the contingency horizon ensures safety. This framework uses a learning cost function $H$ while the primary cost function $J$ appears in additional constraints.
@@
Here, the primary cost function $J$ of the performance horizon is composed of a baseline cost $J_{k}^{B}$, and a deterioration term $\Delta_{k}$ which quantifies the cost incurred for exploration. The deterioration used for exploration $(\Delta_{k})$ is bounded over time by (9c) and at each time step by (9d). A storage variable $Y_{k}:=Y_{k-1}+\bar{\beta}\,\max(J^{+}_{k},0)+\bar{\gamma}-\Delta_{k}$ accumulates cost deterioration, preventing excessive performance loss over time. The term $J^{+}_{k}:=J(x_{k-1},\tilde{x}_{0\mid k-1},\tilde{U}_{k-1})-J^{B}_{k}$, where $\tilde{x}_{0\mid k-1},\tilde{U}_{k-1}$ denote the previous solution of (9), allows exploration when it can potentially improve performance [soloperto_augmenting_2020]. The learning behavior is tuned via the user-defined parameters $\bar{\beta}$, $\bar{\gamma}$, $\beta^{\text{max}}$, and $\gamma^{\text{max}}$, see [soloperto_augmenting_2020] for details. We define the baseline cost $J_{k}^{B}$ as the optimal cost $J(x_{k},\hat{x}^{*}_{0\mid k},\hat{U}^{*}_{0\mid k})$ from the performance horizon of (7), which requires solving both (7) and (9) at each time step. Alternatively, an upper bound on $J^{B}_{k}$ can be computed in different ways (see [soloperto_augmenting_2020]).
@@
$H(x_{k},\hat{x}_{0\mid k},\hat{U}_{k})=-\sum_{i=0}^{N}\Sigma^{d}(\hat{\mathbf{z}}_{i\mid k},\mathcal{I}_{k}).$ (10)
@@ V-B Safety Guarantees
Similar to [soloperto_augmenting_2020], the robust recursive feasiblity of the actively learning dual MPC is ensured by relying on a robust MPC formulation. In contrast to [soloperto_augmenting_2020], we only require constraints to be robustly enforced for the contingency horizon [geurts_baltussen_2025], which can reduce conservatism. The robust recursive feasibility of the contingency MPC (7) is then ensured by using a robust MPC in the contingency horizon [geurts_baltussen_2025].
@@ VI Numerical Example
We perform a numerical study on a nonlinear mechanical system to compare the discussed MPC algorithms and demonstrate the proposed actively learning dual MPC. Specifically, we consider a mass–spring–damper system with a nonparametric unknown spring force, shown in Fig. 1, that is described by the continuous-time state-space equation:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\underbrace{\begin{bmatrix}x_{2}\\ -\frac{1}{m}\left(c_{d}\,x_{2}+c_{k}\,x_{1}+u\right)\end{bmatrix}}_{=\,f_{c}(x,u)}+\underbrace{\begin{bmatrix}0\\ -\frac{1}{m}c_{k}(e^{-x_{1}}-1)x_{1}\end{bmatrix}}_{=\,Bg_{c}(x,u)},$
where $x_{1}$ [m] is the displacement, $x_{2}$ [m/s] is the velocity, $m=1$ [kg] is the mass, $c_{d}=1$ [Ns/m] is the damping constant and $c_{k}=0.33$ [N/m] is the spring constant. We consider $v_{k}=0$. The system is subject to state constraints $\mathbb{X}=[-0.1,1.1]\times[-5,5]$, and input constraints $\mathbb{U}=[-5,5]$. The goal is to control the system to a desired setpoint $x_{\text{ref}}^{\top}=\begin{bmatrix}1&0\end{bmatrix}$ and $x_{\text{ref}}^{\top}=\begin{bmatrix}1.095&0\end{bmatrix}$ after $t=5$ s, while actively learning the unknown nonlinear spring force using the proposed dual MPC. The continuous-time system is linearized at the origin and discretized using the forward Euler method with a sampling time of $0.1$ seconds, yielding the nominal discrete-time model $f$. The input is applied with a zero-order hold and is kept constant over the sampling period. The primary cost function is a standard tracking cost
@@
VI-A MPC Implementation and Properties
We consider the primary cost function (11) with weighting matrices $Q=\text{diag}(20,1)$ and $R=1$ with $P$ from the LQR corresponding to $Q$ and $R$, and $\lambda=10^{-3}$. We consider a prediction horizon of $N=20$. For the GP regression, we use $x_{1}$ as the regression feature with a squared exponential kernel $\mathfrak{K}$ with a signal variance $\sigma_{f}^{2}=0.33$, noise variance $\sigma_{v}^{2}=0$ and length scale $l=0.3$ [kocijan_modelling_2016]. We use $M=4$ inducing points. The unmodeled spring force only affects the velocity dynamics, hence, $B=[0,1]^{\top}$. We consider $\kappa\equiv 0$ and $w_{k}\in\mathbb{W}$ for all $k\in\mathbb{N}$ with $\mathbb{W}=\{0\}\times[-0.33,0.33]$. The parameters of the active learning framework are set to $\bar{\beta}=1$, $\bar{\gamma}=0$, $\beta^{\text{max}}=0$, and $\gamma^{\text{max}}=\infty$, so that the degree of excitation decays over time as the tracking cost decreases. The MPC parameters are constant for all MPC formulations.
The RMPC (3) robustly accounts for all possible realizations of the unmodeled dynamics and enforces robustly tightened constraints with an RCI terminal set, computed using the LQR, to ensure recursive feasibility. The GP-MPCs leverage the online collected measurements to identify and predict the unmodeled dynamics and use adaptive constraint tightening based on the anticipated state covariance. Furthermore, the GP-MPCs use the RMPC (3) in the contingency horizon. The passive GP-MPC [geurts_baltussen_2025] only leverages the cautious property of the dual effect. The proposed dual GP-MPC (9) uses the learning objective (10) to leverage both the cautious and explorative properties of the dual effect (Section III-A). Finally, we evaluate an implementation of the active learning framework by [soloperto_augmenting_2020] that does not use a secondary horizon and relies on robustly tightened constraints from (3), and, therefore, lacks the cautious property of the dual effect.
@@ VI-B Results
The results of the simulation study are shown in Fig. 2. The RMPC considers the worst-case disturbance caused by the uncertain spring force. Initially, the RMPC overshoots and has a large steady-state error of 2.74%. After $t=5$ s, the improved tracking is largely incidental. As the nominal model is unaware of the reduced spring stiffness, the controller tends to overcompensate, and the system happens to settle near the constraint, yielding a small steady-state error of 0.48%, at $t=10$ s. The actively learning GP-MPC with a single horizon [soloperto_augmenting_2020] actively explores the state space and later exploits the learned model. Although the learned model is aware of the reduced stiffness, it is still constrained to the robust constraints that were designed for the worst-case disturbance. Consequently, the MPC cannot fully exploit the learned model, leading to conservative behavior and steady-state errors of 3.16% and 3.53%, at $t=5,10$ s, respectively.
The passively learning GP-MPC with the contingency horizon [geurts_baltussen_2025] is able to successfully control the system to the desired setpoint while learning and adapting to the nonparametric unmodeled dynamics. As the two horizons branch after the first input, the contingency horizon ensures that the RCI set is reachable at all times, while the performance horizon can exploit the learned model to reduce conservatism and improve performance, significantly reducing the steady-state errors to 0.17% and 0.13%, at $t=5,10$ s, respectively. Lastly, the proposed actively learning GP-MPC can safely excite the system to learn the unmodeled dynamics, and exploit the learned model to obtain similar performance (0.19% and 0.13%). The active GP-MPC (9) has an average solve time of $0.295$ seconds on a standard laptop PC using MATLAB, CasADi [andersson_casadi_2019] and IPOPT [wachter_implementation_2006], demonstrating the practical computational efficiency of the proposed approach.
@@ VII Conclusion
In this work, we proposed a Gaussian process (GP)-based dual MPC that leverages the dual effect to actively learn nonparametric uncertainties. We extended the active learning framework of [soloperto_augmenting_2020] by including the dual effect and adopted a multi-horizon contingency MPC formulation [geurts_baltussen_2025] to ensure robust constraint satisfaction. The proposed dual MPC features both caution and exploration, the two properties of dual control, enabling safe active learning and control near constraints, and it was demonstrated on a nonlinear system.
Although a simple problem was considered for the sake of analysis, we expect that the actively learning dual MPC will outperform the passive variant in more complex problems that require exploration to learn the unknown dynamics. Furthermore, the results show that the contingency framework can reduce conservatism compared to a single-horizon MPC. Future work will focus on the performance analysis of both the actively and passively learning dual MPC and explore the use of robust adaptive MPC to further reduce conservatism.